if the degree measure of the smallest angle of the triangle is one-half of the degree measure of the second largest angle and one-third of the degree measure of the largest angle, then what is the degree measure of each angle.
Answer by meghanjax13(2)

the largest angle > the second largest angle
1/3 of the largest angle = 1/2 the second largest angle
So the if largest angle is 90, and 1/3 of 90 is 30, then the smallest angle would be 30.
And because 1/3 of the largest angle = 1/2 of the second largest angle, the second largest angle has to be 60.
Largest angle = 90
second largest angle = 60
smallest angle = 30
Hope I could help.